OpenStudy (anonymous):

show that when \(n=(a+b)=a^2-b^2\) then n is odd .

ganeshie8 (ganeshie8):

@mathmath333

Nnesha (nnesha):

(a+b) is not equal to a^2 - b^2 ^.^ >,< ^.~

OpenStudy (anonymous):

^_^ cute

ganeshie8 (ganeshie8):

it can

ganeshie8 (ganeshie8):

for example : 2 + 1 = 2^2 - 1^2

Nnesha (nnesha):

yeah but........... ~.^ a^2 - b^2 = (a+b) this is right but its also (a -b) :D:D

Nnesha (nnesha):

hmmmmm bye

OpenStudy (anonymous):

this observation makes u solve it @Nnesha

ganeshie8 (ganeshie8):

yes i see now

ganeshie8 (ganeshie8):

thats a good starting point

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{a^2-b^2=a+b\hspace{.33em}\\~\\ (a+b)(a-b)=a+b \hspace{.33em}\\~\\ (a-b)=1 \hspace{.33em}\\~\\ a=b+1 \hspace{.33em}\\~\\ n=a+b \hspace{.33em}\\~\\ n=b+1+b \hspace{.33em}\\~\\ n=2b+1 \hspace{.33em}\\~\\ n(mod~2)=(2b+1)(mod~2)\hspace{.33em}\\~\\ n(mod~2)\equiv 1\hspace{.33em}\\~\\ }\end{align}\) n is odd

OpenStudy (anonymous):

:D yes yes

OpenStudy (mathmath333):

thnks thnks :)

OpenStudy (anonymous):

np np ;)

ganeshie8 (ganeshie8):

nice nice

Nnesha (nnesha):

okay i'm back :D

Nnesha (nnesha):

hmmm O_O that mod thingy :O what is mod ??? mod ganeshie8 ??

Nnesha (nnesha):

@nnesha

OpenStudy (mathmath333):

@Marki

OpenStudy (anonymous):

what mod thingy ? mod in the question is modules :) or residues

Nnesha (nnesha):

:O:P okay

Nnesha (nnesha):

a2−b2=a+b (a+b)(a−b)=a+b (a−b)=1 a=b+1 n=a+b n=b+1+b n=2b+1 oaky so i got this one what you did after this n(mod 2) ????

OpenStudy (anonymous):

n=2b+1 means n is odd :)

OpenStudy (anonymous):

\(\Huge \color{red}{\heartsuit}\)

Nnesha (nnesha):

hmmmm marki law of love with math lololol

OpenStudy (anonymous):

^_^

Nnesha (nnesha):

n(mod 2)=(2b+1)(mod 2) n(mod 2)≡1 <---------- so how he./she got 1 here ?? :O

ganeshie8 (ganeshie8):

Any integer of form \(2a + 1\) is odd, where \(a\) is some integer

ganeshie8 (ganeshie8):

Any integer of form \(2a\) is even, wher e \(a\) is some integer

OpenStudy (anonymous):

yeah mod means residues , when u divide 2b+1 over 1 would left , right ?

ganeshie8 (ganeshie8):

those are the definitions of odd and even numbers

OpenStudy (anonymous):